# GGMotion: Group Graph Dynamics-Kinematics Networks for Human Motion Prediction

### v1

@@ I Introduction
Human motion prediction refers to predicting future human skeletal motion poses based on observed motion sequences, which is an important topic in computer vision and plays a significant role in practical applications such as autonomous driving [1, 2], human-computer interaction [3, 4], and animation production [5]. The human skeleton resembles a rigid body structure composed of multiple joints [6], and 3D human motion encompasses submotions of different body parts. Additionally, human motion is a physical process with complex spatio-temporal dependencies in 3D geometric space, which makes it challenging to model explicitly by deterministic algorithms.
Deep learning has emerged as the predominant approach to human motion prediction tasks. Due to Recurrent Neural Networks (RNN) [7, 8, 9, 10, 11] struggling to capture spatial information, existing studies predominantly focus on Transformer [12, 13, 14], and Graph Convolution Networks (GCN) [15, 16, 17, 18, 19, 20, 21] to model the spatial structure of human poses. The iterative prediction paradigm [22, 23] in a chain-like dependency tends to accumulate errors along the chain, degrading the accuracy of the extremity joint predictions. Meanwhile, [24, 25] have attempted to split the body into upper and lower parts to capture the motion characteristics of different body parts. Further, [26] collaboratively models joints with similar dynamic information to optimize implicit correlations. Despite achieving promising results on benchmark datasets, these approaches remain at the human topology modeling level without delving into the inherent dynamics and kinematics constraints of human motion, for example, joints perform rigid-body motion around their parent joints, which makes them prone to generating physically implausible poses. On the other hand, since human motion occurs in 3D geometric space, it is essential to maintain geometric equivariance under Euclidean transformations (i.e., translation, rotation, and reflection) to achieve better generalization with fewer parameters, as demonstrated in [27, 28].
@@
To incorporate dynamics and kinematics into the network, we improve the equivariant message passing function [29] by integrating the self-attention mechanism of the Transformer. Specifically, we replace the conventional softmax-based attention computation with a covariance matrix-based operation to preserve the equivariance of the MLP. Instead of [24, 25], we divide the human skeleton into multiple groups specifically for the dynamics and kinematics modeling shown in Fig. 1. Compared to full-body modeling, the grouping strategy is more beneficial for learning the differences between various body parts. Based on the equivariant MLP, the inter-group and intra-group interaction modules are designed to capture motion force dependencies at different scales. Finally, rigid-body dynamics and kinematics propagation are employed in each group to update the joint position features in parallel, contributing to more accurate and physically plausible motion predictions. An auxiliary loss based on joint length is introduced to impose prior constraints on human motion.
@@ II-A Human Motion Prediction
The early research on human motion prediction mainly focuses on traditional probabilistic models [30, 31, 32], which directly impose prior knowledge about human motion. Nevertheless, these methods struggle to generate continuous and reasonable human poses. Subsequently, the field of human motion prediction entered an era driven by deep learning. Since human motion sequences exhibit similarities to time series data, Recurrent Neural Networks (RNN) [7, 8, 33, 34, 9, 10, 35, 11, 22] are employed to capture temporal dependencies. To address the inherent cumulative error problem in RNN-based networks, researchers have employed various strategies, e.g., residual structures [11, 10], generative adversarial training [34], and error adaptation mechanisms [7]. Then the spatio-temporal characteristics of human motion motivated researchers to employ sequence-to-sequence convolutional neural networks [36, 37, 38] and attention mechanisms [13, 12, 39, 14] for modeling spatio-temporal dependencies, which predict sequences from sequences and significantly mitigate the cumulative error between predicted frames. Dai et al. [40] introduced manually computed kinematic features as an additional input. Nevertheless, the reliance on handcrafted features constrains the ability to learn robust and expressive motion representations.
@@ III-C Equivariant MLP
$\phi_{eq}(\mathbf{Z}^{n}):=\mathbf{Z}_{\mathrm{v}}^{n}\mathrm{MLP}(\frac{\boldsymbol{\Sigma}^{n}}{||\boldsymbol{\Sigma}^{n}||_{2,row}}),\quad\boldsymbol{\Sigma}^{n}=\mathbf{Z}_{\mathrm{q}}^{n\top}\mathbf{Z}_{\mathrm{k}}^{n},$
@@
The input $\mathbf{Z}$ is mapped to $\mathbf{Z}_{\mathrm{q}}\in\mathbb{R}^{D\times C}$,$\mathbf{Z}_{\mathbf{k}}\in\mathbb{R}^{D\times C}$,$\mathbf{Z}_{\mathbf{v}}\in\mathbb{R}^{D\times C}$ through three linear layers $\mathbf{W}_{\mathrm{q}}\in\mathbb{R}^{D\times C}$,$\mathbf{W}_{\mathrm{k}}\in\mathbb{R}^{D\times C}$,$\mathbf{W}_{\mathrm{v}}\in\mathbb{R}^{D\times C}$. After that, the covariance matrix $\boldsymbol{\Sigma}^{n}$ of the $n$ variables was nonlinearly transformed in MLP, and $\mathbf{Z}_{\mathbf{v}}$ is used to restore the data to the $D$-dimensional coordinate space through matrix multiplication. The final output is produced by a linear layer. Since the rotation (or reflection) matrix $\mathrm{R}$ satisfies $\mathrm{R}^{\top}\mathrm{R}=\mathrm{I}$, after subtracting the centroid of the pose sequence to eliminate the global translation vector, we can easily conclude that $\phi_{eq}(\cdot)$ exhibits equivariance.
@@ III-D Network Architecture
$\mathbf{e}_{ij}^{l}=\mathbf{\beta}_{i}^{l}\cdot\phi_{e}^{l}(||\mathbf{X}_{i}-\mathbf{X}_{j}||_{2,col}),$ (2)
@@
where $\boldsymbol{f}_{i}^{l}\in\mathbb{R}^{D\times C}$ is the force acting on the $i^{th}$ joint, $\sum_{n\in\mathcal{G}_{s}}\boldsymbol{f}_{n}$ is the resultant force in the group $\mathcal{G}_{s}$, [,][$\cdot$,$\cdot$][ ⋅ , ⋅ ] is concatenation. The output of $\phi_{eq}(\cdot)$ is added through a skip connection to the corresponding force.
@@ III-E Loss Function
TABLE I: Comparisons of short-term prediction on Human3.6M. Results at 80ms, 160ms, 320ms and 400ms in the future are shown. All samples are tested for each action. The best results are highlighted in bold, and the second best result is underline.
@@
IV Experiment
IV-A Datasets
IV-B Settings
IV-C Baselines
We compare our method with the state-of-the-art approaches on the benchmark datasets. The MPJPE is evaluated on the entire test set in a 3D coordinate space to provide a more credible and comprehensive reflection of model performance. The compared methods include Res-sup [11], Traj-GCN [19], DMGNN [15], MSRGCN [17], STSGCN [18], siMLPe [41], PGBIG [20], SPGSN [24], EqMotion [28], DPnet [26], and KSOF [21]. Res-sup [11] is a classical RNN-based method that enhances performance through residual connections. DMGNN [15] and MSRGCN [17] abstract the body topology at multiple scales to capture richer spatial information. EqMotion [28] employs an equivariant graph neural network with invariant interaction reasoning to learn the motion features of each node. In addition to GCN, STSGCN [18] and KSOF [21] incorporate the Temporal Convolutional Network (TCN) to capture temporal dependencies. For a fair comparison, we either use their publicly available pre-trained models or re-train them with the default hyperparameters.
@@ IV-C Baselines
millisecond 80 160 320 400 560 1000
Res-sup. [11] 24.0 43.0 74.5 87.2 105.5 136.3
Traj-GCN [19] 9.3 17.1 33.0 40.9 55.8 86.2
DMGNN [15] 13.6 24.1 47.0 58.8 77.4 112.6
MSR-GCN [17] 8.1 15.2 30.6 38.6 53.7 83.0
STSGCN [18] 25.3 27.9 41.8 59.2 65.0 92.1
PGBIG [20] 7.6 14.3 29.0 36.6 50.9 80.1
SPGSN [24] 8.3 14.8 28.6 37.0 51.2 77.8
@@
IV-D Results
We provide the qualitative and quantitative results of our approach in both short-term (i.e., $\leq 500\,\mathrm{ms}$) and long-term (i.e., $>500\,\mathrm{ms}$) prediction.
Results on Human3.6M. We present qualitative comparisons of short-term motion predictions on Human3.6M with several existing methods in Table I. In short-term predictions, our method achieves leading performance in most cases and outperforms baseline methods by a large margin on average. In long-term predictions, our method encounters challenges arising from greater temporal uncertainty and high motion frequency. Although not utilizing temporal modeling methods such as TCN [18, 21] or DCT [19, 17, 20, 24], which are known to benefit sequence prediction, our method still delivers competitive prediction results.
Figure 5: Visualization of predicted poses of different methods on Human3.6M.
@@ IV-D Results
Method MPJPE (mm) RunTime (ms) ModelSize (M)
@@
Results on CMU MoCap and 3DPW. The same comparisons on CMU MoCap and 3DPW are presented in Table III and Table IV respectively. Compared to Human3.6M, our method performs better on CMU Mocap and 3DPW, with additional MPJPE reductions of $0.9/3.5mm$ at $400ms$. This is likely attributed to differences in body topology and joint sampling accuracy, where a more realistic joint distribution results in improved prediction accuracy. Consistent improvements across all benchmarks demonstrate the generalizability and reliability of our human topology modeling strategy.
Qualitative results. We present a qualitative example of predicted poses from different methods in Fig. 5. It can be seen that our method produces more accurate short-term predictions that closely match the ground truth. In long-term predictions, while the performance gap between our approach and the baselines is marginal, our method generates more credible and stable human motion poses.
@@ V Ablation Study
Effect of Auxiliary Loss. In Table IX, we evaluate the contribution of different loss terms during training. As shown, incorporating the auxiliary loss $\mathcal{L}_{\text{aux}}$ leads to the best prediction performance at all timestamps.

### v2

# GGMotion: Group Graph Dynamics-Kinematics Network for Human Motion Prediction

@@ I Introduction
Human motion prediction refers to predicting future human skeletal motion poses based on observed motion sequences, which is an important topic in computer vision and plays a significant role in practical applications such as autonomous driving [1, 2], human-computer interaction [3, 4], and animation production [5]. The human skeleton can be regarded as a rigid body system composed of multiple joints, and 3D human motion consists of submotions performed by different body parts. Additionally, human motion is a physical process with complex spatio-temporal dependencies in 3D geometric space, which makes it challenging to model explicitly by deterministic algorithms.
Deep learning has emerged as the predominant approach to human motion prediction tasks. Due to Recurrent Neural Networks (RNN) [6, 7, 8, 9, 10, 11, 12, 13, 14] struggling to capture spatial information, existing studies predominantly focus on Transformer [15, 16, 17, 18, 19, 20, 21], and Graph Convolution Networks (GCN) [22, 23, 24, 25, 26, 27, 28, 29, 30] to model the spatial structure of human poses. The iterative prediction paradigm [13, 31] in a chain-like dependency tends to accumulate errors along the chain, degrading the accuracy of the extremity joint predictions. Meanwhile, [28, 30] have attempted to split the body into upper and lower parts to capture the motion characteristics of different body parts. Further, [32] collaboratively models joints with similar dynamic information to optimize implicit correlations. Despite achieving promising results on benchmark datasets, these approaches remain at the human topology modeling level without delving into the inherent dynamics and kinematics constraints of human motion, for example, joints perform rigid-body motion around their parent joints, which makes them prone to generating physically implausible poses. On the other hand, since human motion occurs in 3D geometric space, it is essential to maintain geometric equivariance under Euclidean transformations (i.e., translation, rotation, and reflection) to achieve better generalization with fewer parameters, as demonstrated in [33, 34].
@@
To incorporate dynamics and kinematics into the network, we improve the equivariant message passing function [35] by integrating the self-attention mechanism of the Transformer [36]. Specifically, we replace the conventional softmax-based attention computation with a covariance matrix-based operation to preserve the equivariance of the MLP. Instead of [28, 30], we divide the human skeleton into multiple groups specifically for the dynamics and kinematics modeling shown in Fig. 1. Compared to full-body modeling, the grouping strategy is more beneficial for learning the differences between various body parts. Based on the equivariant MLP, the inter-group and intra-group interaction modules are designed to capture motion force dependencies at different scales. Finally, rigid-body dynamics and kinematics propagation are employed in each group to update the joint position features in parallel, contributing to more accurate and physically plausible motion predictions. An auxiliary loss based on joint length is introduced to impose prior constraints on human motion.
@@ II-A Human Motion Prediction
The early research on human motion prediction mainly focuses on traditional probabilistic models [37, 38, 39], which directly impose prior knowledge about human motion. Nevertheless, these methods struggle to generate continuous and reasonable human poses. Subsequently, the field of human motion prediction entered an era driven by deep learning. Since human motion sequences exhibit similarities to time series data, Recurrent Neural Networks (RNN) [6, 7, 8, 9, 10, 11, 12, 14, 13] are employed to capture temporal dependencies. To address the inherent cumulative error problem in RNN-based networks, researchers have employed various strategies, e.g., residual structures [12, 10], generative adversarial training [40], and error adaptation mechanisms [6]. The spatiotemporal characteristics of human motion motivated researchers [41, 42, 43] to employ sequence-to-sequence convolutional neural networks (CNN) to capture complex spatiotemporal dependencies, which predict sequences from sequences and significantly mitigate the cumulative error between predicted frames. The attention mechanism [15, 16, 17, 18, 19, 20, 21] is also highly effective in modeling spatiotemporal sequences. Dai et al. [44] introduced manually computed kinematic features as an additional input. However, reliance on hand-crafted features restricts the ability to learn robust and expressive motion representations.
@@ III-C Equivariant MLP
[…]
@@
The input $\mathbf{Z}$ is mapped to $\mathbf{Z}_{\mathrm{q}}\in\mathbb{R}^{D\times C}$$\mathbf{Z}_{\mathbf{k}}\in\mathbb{R}^{D\times C}$$\mathbf{Z}_{\mathbf{v}}\in\mathbb{R}^{D\times C}$ through three linear layers $\mathbf{W}_{\mathrm{q}}\in\mathbb{R}^{D\times C}$$\mathbf{W}_{\mathrm{k}}\in\mathbb{R}^{D\times C}$$\mathbf{W}_{\mathrm{v}}\in\mathbb{R}^{D\times C}$. After that, the covariance matrix $\boldsymbol{\Sigma}^{n}$ of the $n$ variables was nonlinearly transformed in MLP, and $\mathbf{Z}_{\mathbf{v}}$ is used to restore the data to the $D$-dimensional coordinate space through matrix multiplication. The final output is produced by a linear layer. Since the rotation (or reflection) matrix $\mathrm{R}$ satisfies $\mathrm{R}^{\top}\mathrm{R}=\mathrm{I}$, after subtracting the centroid of the pose sequence to eliminate the global translation vector, we can easily conclude that $\phi_{eq}(\cdot)$ exhibits equivariance.
@@ III-D Network Architecture
[…]
@@
where $\boldsymbol{f}_{i}^{l}\in\mathbb{R}^{D\times C}$ is the force acting on the $i^{th}$ joint, $\sum_{n\in\mathcal{G}_{s}}\boldsymbol{f}_{n}$ is the resultant force in the group $\mathcal{G}_{s}$, $\cdot$$\cdot$] is concatenation. The output of $\phi_{eq}(\cdot)$ is added through a skip connection to the corresponding force.
@@ III-E Loss Function
TABLE I: Comparisons of short-term prediction on Human3.6M. Results at 80ms, 160ms, 320ms and 400ms in the future are shown. All samples in the test set are tested for each action. The best results are highlighted in bold, and the second best result is underline.
@@
IV Experiment
IV-A Datasets
IV-B Settings
IV-C Baselines
We compare our method with the state-of-the-art approaches on the benchmark datasets. The MPJPE is evaluated on the entire test set in a 3D coordinate space to provide a more credible and comprehensive reflection of model performance. The compared methods include Res-sup [12], Traj-GCN [26], DMGNN [22], MSRGCN [24], STSGCN [25], siMLPe [45], PGBIG [27], SPGSN [28], EqMotion [34], DPnet [32], and KSOF [29]. Res-sup [12] is a classical RNN-based method that enhances performance through residual connections. DMGNN [22] and MSRGCN [24] abstract the body topology at multiple scales to capture richer spatial information. EqMotion [34] employs an equivariant graph neural network with invariant interaction reasoning to learn the motion features of each node. DPnet [32] proposes a dynamic pattern-based collaborative modeling approach to capture joint dynamic information. In addition to GCN, STSGCN [25] and KSOF [29] incorporate the Temporal Convolutional Network (TCN) to capture temporal dependencies. For a fair comparison, we either use their publicly available pre-trained models or re-train them with the default hyperparameters.
@@ IV-C Baselines
IV-D Results
We provide the qualitative and quantitative results of our approach in both short-term (i.e., $\leq 500\,\mathrm{ms}$) and long-term (i.e., $500\,\mathrm{ms}$) prediction.
Results on Human3.6M. We present qualitative comparisons of short-term motion predictions on Human3.6M with several existing methods in Table I. In short-term predictions, our method achieves leading performance in most cases and outperforms baseline methods by a large margin on average. In long-term predictions, our method encounters challenges arising from greater temporal uncertainty and high motion frequency. Although not utilizing temporal modeling methods such as TCN [25, 29] or DCT [26, 24, 27, 28], which are known to benefit sequence prediction, our method still delivers competitive prediction results, as demonstrated in Table II.
@@ IV-D Results
millisecond 80 160 320 400 1000
Res-sup. [12] 24.0 43.0 74.5 87.2 136.3
Traj-GCN [26] 9.3 17.1 33.0 40.9 86.2
DMGNN [22] 13.6 24.1 47.0 58.8 112.6
MSR-GCN [24] 8.1 15.2 30.6 38.6 83.0
STSGCN [25] 10.8 18.2 31.2 41.2 81.7
PGBIG [27] 7.6 14.3 29.0 36.6 80.1
SPGSN [28] 8.3 14.8 28.6 37.0 77.8
@@
Results on CMU MoCap and 3DPW. The same comparisons on CMU MoCap and 3DPW are presented in Table III and Table IV respectively. Compared to Human3.6M, our method performs better on CMU Mocap and 3DPW, with additional MPJPE reductions of $0.9/3.5mm$ at $400ms$. This is likely attributed to differences in body topology and joint sampling accuracy, where a more realistic joint distribution results in improved prediction accuracy. Relative to the dynamic pattern-based collaborative modeling in DPnet [32], our refined grouping strategy and dynamic-kinematic modeling provide notable improvements. Consistent improvements across all benchmarks demonstrate the generalizability and reliability of our human topology modeling strategy.
Qualitative results. We present a qualitative example of predicted poses from different methods in Fig. 5. It can be seen that our method produces more accurate short-term predictions that closely match the ground truth, especially near the body extremities. In long-term predictions, while the performance gap between our approach and the baselines is marginal, our method generates more credible and stable human motion poses. From the perspective of motion scenarios, our method achieves superior results in both the large-amplitude action ”walking” and the small-amplitude action ”taking photo”.
Figure 5: Visualization of predicted poses of different methods on Human3.6M. The first row presents the ground truth in gray, while the second row displays our predicted results in color. The predicted poses where our method achieves more accurate predictions are highlighted with red boxes.
@@
Method MPJPE ($m$) RunTime (ms) ModelSize (M)
@@ V Ablation Study
Effect of Auxiliary Loss. In Table IX, we evaluate the contribution of different loss terms during training. As shown, incorporating the auxiliary loss $\mathcal{L}_{\text{aux}}$ leads to the best prediction performance at all timestamps. These results demonstrate the crucial role of incorporating the human skeletal length loss term in supervising the training process.